OpenStudy (anonymous):

Two real numbers are selected from the closed interval [0,4] say x and y. What is the probability that the selected numbers sastisfy y^2<=x

OpenStudy (anonymous):

We are choosing points (x,y) in the square [0,4] x[0,4] which has an area of 16

OpenStudy (anonymous):

To satisfy the inequality (x,y) should be chosen under the curve \[ y= \sqrt x \] The area under this curve is \[ \int_0^4 \sqrt x dx=\frac {16}3 \]

OpenStudy (anonymous):

Hence the answer is \[ \frac {\frac {16}3} {16}=\frac 1 3 \]

OpenStudy (anonymous):

awesome!thanks :D